OpenStudy (anonymous):

Solve the system algebraically. 2x + y - 10 = 0 x - y - 5 = 0 What is the value of y? -5/3 0 5

OpenStudy (shinalcantara):

2x + y - 10 = 0 x - y - 5 = 0 or you can have it as 2x + y = 10 -----equation 1 x - y = 5 ----equation 2 in this case, you can do elimination method. since you are to find the value of y. then we would wish to eliminate x multiplying equation 2 by 2 you'll have: x - y = 5 x2 2x - 2y = 10 equation 2' subtract equation 2' from equation 1 2x + y = 10 - 2x - 2y = 10 3y = 0 y=0

OpenStudy (anonymous):

Thanks . :)